. Consider the Schrödinger equation for a particle in the presence of the potential of the form 𝑉(𝑥) = 𝑈(𝑥) + 𝑖𝑊(𝑥) where 𝑈 and 𝑊 are real functions of 𝑥. What form does the conservation equation take?


1
Expert's answer
2021-09-17T09:22:52-0400

let us consider the Schrödinger equation with potential "\ud835\udc49(\ud835\udc65) = \ud835\udc48(\ud835\udc65) + \ud835\udc56\ud835\udc4a(\ud835\udc65)":


"i\hbar \frac{\partial \psi}{\partial t}=-\frac{\hbar^2}{2m}\nabla^2\psi+( 𝑈(𝑥) + 𝑖𝑊(𝑥))\psi\quad (1)"

The conjugate equation

"-i\hbar \frac{\partial \psi*}{\partial t}=-\frac{\hbar^2}{2m}\nabla^2\psi^*+( 𝑈(𝑥) - 𝑖𝑊(𝑥))\psi^*\quad (2)"

After multiplying Eqn 1 by "\psi^*", Eqn 2 by "\psi" and subtracting, we get

"i\hbar\left( \frac{\partial \psi}{\partial t}\psi^*+\frac{\partial \psi^*}{\partial t}\psi\right)="

"-\frac{\hbar^2}{2m}\left(\psi^*\nabla^2\psi-\psi\nabla^2\psi^*\right)+2iW(x)\psi\psi^*"

or

"\frac{\partial \rho}{\partial t}+\nabla {\bf j}-\gamma\rho=0"

Here

"\rho=\psi\psi^*"

"{\bf j}=\frac{\hbar}{2mi}\left(\psi^*\nabla\psi-\psi\nabla\psi^*\right)"

"\gamma=2W(x)/\hbar"